OpenStudy (anonymous):

find the 15th term of (x+2y)^20, where x^20 is labeled as the 0th term.

OpenStudy (rajee_sam):

Do you know the formula for binomial expansion?

OpenStudy (anonymous):

yes

OpenStudy (rajee_sam):

so for finding nth term in a binomial expansion you use the following formula \[nC _{r-1}a ^{n-(r-1)}b ^{(r-1)}\]

OpenStudy (anonymous):

Or PASCAL'S TRIANGLE

OpenStudy (rajee_sam):

r the no. of the term you are trying to find and n is the power of the expansion So for this problem n = 20 and r = 15

OpenStudy (rajee_sam):

\[20C _{14} X ^{(20-14)} (2Y)^{14}\]

OpenStudy (rajee_sam):

\[20C _{14} X ^{6} (2Y)^{14}\]

OpenStudy (rajee_sam):

Can you find the value?

OpenStudy (anonymous):

no, i messed up somewhere.

OpenStudy (anonymous):

can you guide me through the next steps?

OpenStudy (rajee_sam):

how do you find\[20C _{14}\] ??

OpenStudy (anonymous):

20!/r!(n-r)!

OpenStudy (anonymous):

20!/14!(6)!

OpenStudy (rajee_sam):

easier way is the difference between 20 and 14 is 6 right. so you have to write it like this\[\frac{ 20 \times 19 \times 18 \times 17 \times 16 \times 15 }{ 1 \times 2 \times 3 \times 4 \times 5 \times 6 }\] Numerator 6 nos. starting from 20 multiplied and denominator 6 nos. starting from 1 multiplied. Now you simplify this fraction

OpenStudy (rajee_sam):

what you wrote will be the same too

OpenStudy (anonymous):

38760?

OpenStudy (rajee_sam):

now you have 20C 14. The nest term x^6 can't do anything to it so leave it as it is. next term (2y) ^14 = 2^14 y ^14. You can find what 2^14 using calculator. Multiply the answer with what you get for 20C14. That will be your numerical part. Then you will have x^6 y^14 next to it.

OpenStudy (rajee_sam):

yes 38760

OpenStudy (rajee_sam):

then find 2^14

OpenStudy (anonymous):

2^14=16384

OpenStudy (anonymous):

38760*16384*x^6*y^14

OpenStudy (rajee_sam):

38760* 16384

OpenStudy (rajee_sam):

multiply the numbers and write as one

OpenStudy (anonymous):

635043840*x^6*y^14

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

thats it?

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

thanks

OpenStudy (rajee_sam):

that will be your 15th term